OpenStudy (anonymous):

Find a function g and a number A such that 4+ integral from A to x g(t) /t^2 dt = 2qt x

OpenStudy (anonymous):

do you mean \[4+\int_A^x{g(t)\over t^2}dt=2qtx\] ?

OpenStudy (anonymous):

2 sqaure root of x

OpenStudy (anonymous):

\[4+ \int\limits_{A}^{x} \frac{ g(t) }{ t^2 } dt = 2 \sqrt{x}\]

OpenStudy (anonymous):

ok. now differentiate the eq. with respect to "x"